### v4

# Wholesale Market Participation of DERA: Competitive DER Aggregation

Abstract
We consider the aggregation of distributed energy resources (DERs) by a profit-seeking aggregator participating directly in wholesale market under distribution network access constraints. We propose a competitive DER aggregator (DERA) model that maximizes the DERA’s profit while ensuring each aggregated customer gains no less surplus and pays no higher energy cost than under the regulated retail tariff. The DERA participates in wholesale electricity market as virtual storage with optimized generation offers and consumption bids derived from our competitive aggregation model. Also derived are DERA’s bid curves for the distribution network access and DERA’s profitability when competing with the regulated retail tariff. We show that, with the same distribution network access, the proposed DERA’s wholesale market participation achieves the same welfare-maximizing outcome as when its customers participate directly in the wholesale market. Numerical studies compare the proposed DERA with existing methods in terms of customer surplus and DERA profit. We empirically evaluate how many DERAs can survive in the competition at long-run equilibrium, and assess the impacts of DER adoption levels and distribution network access on short-run market outcomes.
Keywords
distributed energy resources and aggregation, behind-the-meter distributed generation, demand-side management, net energy metering, competitive wholesale market.
I Introduction
We address open problems in the direct participation of distributed energy resource aggregators (DERAs) in the wholesale electricity market operated by regional transmission organizations and independent system operators (RTOs/ISOs), under the general framework in FERC order $2$ [2]. We focus on the aggregation strategy of a profit-seeking DERA, whose industrial, commercial, and residential customers have competing service providers, such as their incumbent regulated utilities. The central theme of this work is to develop profitable and competitive aggregation strategies to attract and retain customers. By competitive aggregation, we mean that the benefits of the DERA customers must be no less than those offered by electricity provider benchmarks. An example of such a benchmark is the incumbent utility or a community choice aggregator (CCA) adopting net energy metering (NEM) policies, offering strong incentives to prosumers with behind-the-meter (BTM) DERs [3, 4, 5]. A major barrier to DERA’s entrance to direct wholesale market participation is having an aggregation strategy and a participation model to make DER aggregation a profitable venture[6].
The technical challenge of designing a profitable and competitive DER aggregation is twofold. First, the DERA plays a dual role in the aggregation process: an energy supplier to its customers in retail market and a producer/demand in wholesale market. Its aggregation must consider retail competition, distribution network access limits, and its overall revenue adequacy. To this end, a DERA needs to derive profit-maximizing bids/offers from its competitive aggregation.
Second, competitive aggregation requires the DERA to offer more attractive pricing to its customers than the regulated tariff and shield them from the volatility of wholesale market prices. Examples of unstable pricing are two-part tariffs from Griddy [7] and Amber [8] defined by the wholesale spot price and a connection charge. Although Griddy’s aggregation offered competitive pricing compared to regulated utility tariffs, its customers experienced a 100-fold price surge during the extreme winter storm Uri in 2021.
I-A Related Work
FERC Order $2$ removes regulatory barriers for DERAs to participate in wholesale capacity, energy, and ancillary service markets. Prior studies have demonstrated DERs’ capability to provide ancillary services both empirically [9, 10] and analytically [11]. In this paper, we focus on DERA participation in energy market, where aggregators submit quantity bids [12] or price-quantity bids [13, 14], and the ISO clears the market and issues dispatch signals. Our analysis centers on market efficiency, DERA profitability, and how aggregators directly control DERs [15] to follow dispatch signals. While this study focuses on energy markets, the proposed framework is also extensible to capacity and ancillary service markets.
The growing literature on DER aggregation and wholesale market participation models broadly falls into two categories. One is through a retail market design operated by a distribution system operator (DSO) [16, 17, 18], an aggregation/sharing platform [19, 20, 21], or an energy coalition such as CCA [22, 23, 24]. For the most part, these works do not consider a profit-maximizing DERA’s active participation in the wholesale market. In particular, in [16, 19], the DSO or an aggregation platform participates in the wholesale markets with the aggregated net demand (or possibly net production), treating the wholesale market as a balancing resource.
Our approach belongs to the second category of DER aggregations, where profit-seeking DERAs aggregate both generation and flexible demand resources, participating directly in the wholesale market with bid/offer curves. To ensure secure distribution network operation, DERA obeys the allocated distribution network access limits[25] (or operating envelopes [14]), rather than considering the computationally expensive network power flow constraints. Within the framework of FERC order $2$, this type of DER aggregation has the potential to improve the overall system efficiency and reliability.
Although the notion of competitive DER aggregation has not been formally defined, two prior works have developed competitive aggregation solutions in [22, 12]. In [22], Chakraborty et al. consider DER aggregation by a CCA, where the authors provide an allocation rule that offers its customers competitive benefits with respect to the regulated utility.
Most relevant to our work is the DERA’s wholesale market participation method developed by Gao et al. [12] where the authors consider a profit-seeking DERA aggregating BTM distributed generations (DGs) and offering its aggregated generation resources to the wholesale market. In particular, their approach achieves a social surplus equal to that achievable by customers’ direct participation in the competitive wholesale market. In other words, their approach achieves the highest economic efficiency for aggregating DGs. A significant difference between [12] and this paper is that we formulate a general competitive aggregation problem that includes the regulated utility. In achieving DERA’s profit maximization, our DER aggregation and market participation models are also different from [12].
The approach proposed in [12] follows the earlier work in [13] where a Stackelberg game-theoretic model is used. Both approaches assume that the DERA elicits prosumer participation with an optimized (one-part or two-part) tariff, and the prosumer responds with its quantity to be aggregated by the DERA. The real-time wholesale market price is reflected by the variable price in [7, 8, 12]. Such a variable price conveys low but volatile wholesale prices directly to customers. To protect customers from price spikes in real-time wholesale price, methods like price caps [26] were proposed.
I-B Summary of Results, Contributions, and Limitations
In this paper, we considerably extend the DERA aggregation model in [1] to one that aggregates customers across multiple locations in distribution networks and incorporates security constraints on network injection and withdrawal limits. We further investigate the competitive aggregation impact on market efficiency, price stability, and long-run equilibrium.
First, we propose a DER aggregation approach based on a constrained convex optimization that maximizes DERA surplus while providing higher customer surpluses than those offered by a competing aggregation model. In particular, we are interested in aggregation schemes that are competitive with the regulated utility rates such as various versions of regulated NEM rates, [^1] NEM analyzed in [27] is an inclusive parametric tariff design that captures key features of the existing and proposed NEM tariffs. with which a customer can make cost-benefit comparisons in her decision to become a customer of the DERA. We show that such a competitive DER aggregation, despite that the aggregation involving real-time wholesale locational marginal price (LMP), has an energy cost no greater than the regulated NEM tariff. This implies that the proposed DER aggregation mechanism ensures price stability regardless of the volatility of the wholesale market LMP, a property missing in Griddy’s pricing model [7]. Meanwhile, we establish the profitability of DERA when competing with NEM.
Second, we propose a virtual storage model for DERA’s wholesale market participation compatible with the practical continuous storage facility participation considered by ISOs under FERC order 841 [28, 29]. The DERA bidding curve is derived from the closed-form solution of the proposed DERA model. While the aggregation optimization explicitly involves wholesale market LMP, the virtual storage bidding curves do not require forecasting of LMP. We show that the proposed DERA wholesale market participation results in market efficiency equal to what is achievable when DERA’s customers participate directly in the wholesale market.
Finally, we derive the benefit function of DERA over distribution network injection and withdrawal access limits. DERAs compete in the distribution network access auction proposed by [25] to acquire network access and we empirically evaluate the number of surviving DERAs in the long-run competitive equilibrium. We also present a set of numerical results, comparing the surplus distribution of the proposed competitive aggregation solution with those of various alternatives, including the regulated utility. Among significant insights gained are the higher social surplus, customer surplus, and DERA surplus achievable in the proposed competitive DERA model, when compared to other alternatives.
A few remarks are warranted regarding the scope and limitation of this paper. First, the losses in distribution systems are not considered. Second, the contingency cases where DSO rejects cleared bids and offers from DERA for reliability concerns [2] are neglected. Under the access limit allocation framework proposed in [25], reliability concerns of DER aggregation are already satisfied under normal operating conditions. Lastly, although the proposed competitive aggregation offers higher benefits to DERA customers, it does so with a non-uniform payment, which might raise equity concerns.
TABLE I: Major symbols
${\bm{d}}$: consumption bundle of aggregated customers.
$\overline{{\bm{d}}},\underline{{\bm{d}}}$: consumption bundle’s upper and lower limits.
$\overline{{\bm{C}}},\underline{{\bm{C}}}$: distribution network injection and withdrawal access limits.
$g,G$: BTM single and aggregated DG.
${\cal K}$: competitiveness constant for benchmark prosumer surplus.
$N$: total number of prosumers.
$M$: total number of points of aggregation (PoAs).
${\cal N}_{m}$: set of aggregated customers under the $m$-th PoA.
$\omega$: payment function of the aggregated customer.
$\pi^{+},\pi^{-},\pi^{0}$: import rate, export rate, and fixed charges of NEM.
$\pi$: wholesale locational marginal price (LMP).
$Q$: aggregated net injection quantity of DERA.
$S_{\mbox{\tiny DERA}}$: total surpluses of DERA and its aggregated prosumers.
$S_{\mbox{\tiny NEM}}$: prosumers surplus under tariff NEM.
$\mathscr{S}(\cdot)$: aggregated supply function.
$U$: prosumer utility function for energy consumption.
$V$: prosumer marginal utility function.
I-C Paper Organization and Notations
In Sec. II, we summarize the DER aggregation model and its main interactions. The problem of competitive DER aggregation is formulated in Sec. III where we derive the optimal aggregation solution. Sec. IV and Sec. V consider DERA’s wholesale market participation and its bidding strategies in the distribution network access auction, respectively. Numerical simulations are presented in Sec. VI. Proofs are relegated in the appendix.
A list of major designated symbols is shown in Table I. The notations used here are standard. We use boldface letters for column vectors as in ${\bm{x}}=(x_{1},\ldots,x_{n})$. In particular, $\bm{1}$ is a column vector of all ones. The indicator function is denoted by $\mathbbm{1}\{x_{n}\leq y_{n}\}$, which equals 1 if $x_{n}\leq y_{n}$, and 0 otherwise. ${\bm{x}}\preceq{\bm{y}}$ means $x_{n}\leq y_{n},\forall n$. $\mathbb{R}_{+}$ represents the set of all nonnegative real numbers. $[x]$ represents the set of integers from 1 to $x$, i.e., $[x]:=\{1,\ldots,x\}$.
II DER Aggregation Model
A DERA aggregates flexible resources from its customers and coordinates with the DSO for power delivery to the wholesale market operated by ISO/RTO. Following the DERA interaction model proposed in [25], we focus on the DERA-DSO-ISO/RTO interfaces (a)–(c), as shown in Fig. 1. Since a DERA uses DSO’s physical infrastructure for power delivery between its customers and the wholesale market, it is essential to delineate the financial and physical interactions at these interfaces. Below, we describe the three interfaces (a)–(c).
Figure 1: DERA model’s physical and financial interactions. The red arrows show the bidirectional power flow, the green for the financial transactions, and the blue for direct control signals.
DERA and its customers at interface (a): We assume the DERA aggregates resources in a retail market from residential, commercial, and industrial customers, who have the option of being served by a regulated utility. Under a single-bill payment model, each customer settles both consumption charges and production compensation with the DERA. The DERA deploys an energy management system with direct controls [^2] The direct control in [15] is implemented through cloud-based platforms. of the customer’s BTM generation and flexible demand resources, such as rooftop PV, heat pumps, water heaters, and EV chargers. To remain competitive with the incumbent utility, the DERA must offer more attractive aggregation benefits; otherwise, customers will revert to the utility. Accordingly, the DERA optimizes the customer’s BTM resources and provides the customer with a cost-benefit comparison with the NEM benchmark offered by the utility. See Sec. III for details.
DERA and RTO/ISO at interface (b): We focus on DERA’s participation in the energy market based on a virtual storage model compatible with the continuous storage facility participation model [28]. See Sec. IV for the construction of bid/offer curves. To this end, the DERA submits offer/bid curves or self-scheduled quantity bids. The DERA may participate in both the day-ahead and real-time markets, although here we focus only on the real-time energy market participation. The DERA may also deploy its own DG and storage capabilities to mitigate aggregation uncertainties.
DERA and DSO at interface (c): We consider the DERA-DSO coordination model in [25], where DERA acquires access limits at distribution network buses operated by DSO. DERA’s willingness to pay for network access is explained in Sec.V. Specifically, DERA secures injection and withdrawal access either through an access limit auction or a bilateral contract with the DSO. During real-time operation, DERA must aggregate DER from its customers in such a way that abides by the injection and withdrawal constraints set by the allocated access limits. That way, DERA’s aggregation has no effect on the operational reliability of DSO under nominal operating conditions, avoiding DSO intervention on ISO dispatch of DERA’s aggregation.
In summary, these three interactions at (a)–(c) establish our core framework for a DER aggregation model that is competitive, profit-making, and grid-aware.
III Competitive DER Aggregation
This section formulates the optimal competitive aggregation and analyzes the properties of the optimal solution when competing with the incumbent utility’s NEM. Our DER aggregation is built on the deregulated retail market. For example, in Texas and New York, [^3] For reference, we provide links to publicly available listings of retail energy providers in Texas and New York. customers can choose their electricity suppliers based on electricity rate and services. We consider prosumers owning all energy consumption and DG devices. After joining a DERA, prosumers grant device access to DERA for measurements and direct controls.
III-A Closed-Form Solution for Competitive DER Aggregation
We consider a DERA aggregating customers over multiple points of aggregation (PoAs) in the distribution network. [^4] For simplicity, we illustrate the single time interval aggregation model here and apply it to the multi-interval aggregation empirically in Sec. VI-E. We define PoAs [^5] A diagram illustrating PoA is in Fig. 2 of [25]. as the main buses with higher voltages in the distribution network, which can be recognized with main substation information [30]. With the DERA-DSO coordination method in [25], [^6] Details about how the DERA coordinates with DSO to get distribution network injection and withdrawal accesses are explained in Sec. V. DERA receives injection and withdrawal access limits at all PoAs, represented respectively by
$\overline{{\bm{C}}}:=(\overline{C}_{m},m\in[M]),~{}~{}\underline{{\bm{C}}}:=(\underline{C}_{m},m\in[M]),$
where $\overline{{\bm{C}}},\underline{{\bm{C}}}\in\mathbb{R}_{+}^{M}$ and $M$ denotes the total number of PoAs. Thus, in the real-time operation, DERA must guarantee that its aggregated power at the $m$-th PoA satisfies
$-\underline{C}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}(g_{n}-\mathbf{1}^{\top}{\bm{d}}_{n})\leq\overline{C}_{m},\forall m\in[M],$ (1)
where $g_{n}\in\mathbb{R}_{+}$ represents the BTM DG output of the $n$-th aggregated customer and ${\cal N}_{m}$ the set of aggregated customers under the $m$-th PoA. Denote $N$ as the total number of DERA customers, and mapping $\rho(n):[N]\rightarrow[M]$ such that $\rho(n)$ gives the index of PoA connecting customer $n$, then
${\cal N}_{m}:=\{n\in[N]~{}|~{}\rho(n)=m\}.$ (2)
${\bm{d}}_{n}\in\mathbb{R}_{+}^{K}$ is the consumption bundle of all customers, i.e.,
${\bm{d}}_{n}:=(d_{nk},k\in[K]),$
where $K$ denotes the total number of energy-consuming devices, including lamps, air-conditioners, washers/dryers, and heat pumps, for each customer $n\in[N]$. Customers set exogenous parameters $\underline{{\bm{d}}}_{n},\bar{{\bm{d}}}_{n}\in\mathbb{R}_{+}^{K}$ as the minimum and maximum energy consumption limits of each device, i.e.,
$\underline{{\bm{d}}}_{n}\preceq{\bm{d}}_{n}\preceq\bar{{\bm{d}}}_{n},\forall n\in[N].$ (3)
Feasibility of the DER aggregation requires that the distribution network access limits (1) and consumption limits (3) have a non-empty intersection at all times. Thus, we assume DERA acquires enough injection and withdrawal accesses such that
$\sum\limits_{n=1}^{{\cal N}_{m}}\sum\limits_{k=1}^{K}\underline{d}_{nk}-\underline{C}_{m}\leq\sum\limits_{n=1}^{{\cal N}_{m}}g_{n}\leq\sum\limits_{n=1}^{{\cal N}_{m}}\sum\limits_{k=1}^{K}\overline{d}_{nk}+\overline{C}_{m}.$ (4)
To attain customers in the energy aggregation, DERA adopts the ${\cal K}$-competitive constraint in (5) to ensure that the surplus of each prosumer under aggregation is higher than the benchmark surplus ${\cal K}_{n}$ (e.g. surplus under the incumbent provider), i.e.,
$U_{n}({\bm{d}}_{n})-\omega_{n}\geq{\cal K}_{n},\forall n\in[N].$ (5)
$U_{n}({\bm{d}}_{n})$ is the $n$-th customer’s utility of consuming ${\bm{d}}_{n}$. We assume the utility function is concave, nonnegative, nondecreasing, continuously differentiable, additive (i.e., across the $K$ devices $U({\bm{d}})=\sum_{k=1}^{K}U_{k}(d_{k})$). Here, the utility function is given; in practice, utility functions can be computed by parametric [27] or nonparametric [31] methods.
${\cal K}$-competitive constraint is the criterion for a rational customer, seeking surplus maximization, to join a DERA. Otherwise, a rational prosumer has the incentive to leave DERA and switch to the benchmark service provider for a higher customer surplus. More details about the benchmark prosumer surplus ${\cal K}_{n}$ are explained in Sec. III-B.
[…]
${\bm{D}}:=(\bm{d}_{n},n\in[N]),~{}~{}\hbox{\boldmath$\omega$\unboldmath}:=(\omega_{n},n\in[N])$
from the following convex profit maximization
[…]
The optimal value $\Pi(\overline{{\bm{C}}},\underline{{\bm{C}}})$ is DERA’s profit given the distribution network access. In the objective function, DERA seeks profit maximization from both the aggregated customers’ payments and the revenue from the wholesale market. Without loss of generality, all PoAs are under the same point of interconnection, facing a common LMP $\pi\in\mathbb{R}$.
Under the above assumptions for feasibility in (4) and the utility function, Theorem 1 establishes a closed-form solution of (6) parameterized by $\pi$.
Theorem 1 (Optimal DERA scheduling and payment).
Given the wholesale market LMP $\pi$, the optimal consumption bundle ${\bm{d}}^{*}_{n}=(d^{*}_{nk})$ of prosumer $n$ and its payment $\omega_{n}^{*}$ are given by
$\begin{array}[]{l}d^{*}_{nk}=\begin{cases}h_{nk}(\underline{\xi}_{m}),&\sum\limits_{n=1}^{{\cal N}_{m}}g_{n}\leq\sum\limits_{n=1}^{{\cal N}_{m}}\sum\limits_{k=1}^{K}h_{nk}(\pi)-\underline{C}_{m}\\ h_{nk}(\overline{\xi}_{m}),&\sum\limits_{n=1}^{{\cal N}_{m}}g_{n}\geq\sum\limits_{n=1}^{{\cal N}_{m}}\sum\limits_{k=1}^{K}h_{nk}(\pi)+\overline{C}_{m}\\ h_{nk}(\pi),&\text{otherwise}\end{cases}\end{array}$ (7)
$\begin{array}[]{lrl}\omega^{*}_{n}&=&U_{n}({\bm{d}}_{n}^{*})-{\cal K}_{n},\end{array}$ (8)
where $V(x):=\frac{d}{dx}U(x)$ is the marginal utility function and
$h_{nk}(x):=\max\{\underline{d}_{nk},\min\{V_{nk}^{-1}(x),\overline{d}_{nk}\}\}.$ (9)
$m$ is the index of PoA connecting customer $n$, i.e., $n\in{\cal N}_{m}$ defined by (2). Solve for $\underline{\xi}_{m},\overline{\xi}_{m},\forall m\in[M]$ from
$\sum_{n=1}^{{\cal N}_{m}}\sum_{k=1}^{K}h_{nk}(\underline{\xi}_{m})=\sum_{n=1}^{{\cal N}_{m}}g_{n}+\underline{C}_{m},$ (10)
$\sum_{n=1}^{{\cal N}_{m}}\sum_{k=1}^{K}h_{nk}(\overline{\xi}_{m})=\sum_{n=1}^{{\cal N}_{m}}g_{n}-\overline{C}_{m}.$ (11)
The proof is in Appendix VIII-A, following the convexity and Karush-Kuhn-Tucker (KKT) conditions of (6). This optimal solution has two noteworthy characteristics. First, the optimal consumption in (7) is only a function of the LMP $\pi$ when DERA purchases enough network accesses at the PoA. Note also the difference between the optimal consumption schedule in (7) and those in [13] where the optimally scheduled consumption always depends on the anticipated LMP and forecast of BTM DG. Second, (6) finds a Pareto efficient allocation that maximizes the surplus of the DERA, subject to the constraint that the aggregated customer has the given level of surplus ${\cal K}_{n}$. Similar optimization and the Pareto efficient allocation are also analyzed in [32, P602] for first-degree price discrimination. The payment function $\omega_{n}$ can be realized by a two-part tariff, which is explored in [12], although it unidirectionally aggregates BTM DG. Overall, such a closed-form solution allows DERA to apply simple dispatch and pricing policies when aggregating massive numbers of households over multiple PoAs in the distribution networks.
III-B Properties of DERA Competitive with NEM X
We analyze the profitability of DERA and the energy consumption cost of aggregated prosumers when DERA is competitive with a regulated NEM tariff parameterized by the retail (consumption) rate $\pi^{+}$, the sell (production) rate $\pi^{-}$, and the connection charge $\pi^{0}$. Assume $0\leq\pi^{-}\leq\pi^{+}$ and customers’ surpluses under NEM are nonnegative [27]. Denote the $n$-th prosumer surplus under NEM to be $S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n})$, whose computation depends on the DG generation and network access limits. Explicit formulations are in Sec. VIII-B. DERA sets the benchmark prosumer surplus
${\cal K}_{n}=\zeta S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n}),$ (12)
which is used in (5) with profit ratio $\zeta\geq 1$ to obtain competitive aggregation over the DSO’s NEM-based aggregation with the same network access. [^7] Customers owning DERs switch from NEM X to DERA for higher consumer surplus, granting DERs control to DERA upon joining. In this subsection, the network access limits carry the subscript $n$, which is equivalent to $m$ since we set ${\cal N}_{m}=1$ and $K=1$ for simplicity.
The ${\cal K}$-competitive constraint in (5) has significant implications on pricing stability, despite that the aggregation is based on real-time LMP. Price stability means the price and payment faced by customers cannot go randomly high, for which a counterexample is the real-time LMP. Because the NEM tariff has price stability, achieving a finite customer payment regardless of the wholesale LMP fluctuation, an aggregation mechanism competitive with the NEM tariff must also be stable. The proposition below formalizes this intuition.
Proposition 1 (Average cost of consumption).
The prosumer’s average energy consumption cost under DERA aggregation is no higher than the NEM retail rate, i.e., $\omega_{n}^{*}/d_{n}^{*}\leq\pi^{+}$.
The proof is provided in Appendix VIII-C. Such price stability comes directly from the ${\cal K}$-competitive constraint, which enforces a lower bound for customer surplus and thus naturally limits the maximum customer payment. Note that the two-part pricing of Griddy [7] is not a stable pricing mechanism because the retail rate is tied directly to the real-time LMP.
In the ${\cal K}$-competitive constraint (5) with (12), the profit ratio $\zeta$ controls surplus distribution between DERA and its aggregated prosumers. A larger $\zeta$ rebates more benefits to prosumers and incentivizes prosumers to join DERA, although it increases the deficit risk of DERA. Therefore, the DERA must carefully set $\zeta$ to balance profitability and competitiveness. In Proposition 2, we establish that the DERA can achieve nonnegative expected profit by choosing an appropriate $\zeta$, assuming that BTM DG generation ${\bm{g}}$ and the LMP $\pi$ are independent random variables in a competitive market.
Proposition 2 (Profitability of DERA).
If $\pi^{-}\leq\mathbb{E}[\pi]\leq\pi^{+}$, then there exists a profit ratio $\zeta\geq 1$ such that the DERA’s expected profit is nonnegative.
The proof is provided in Appendix VIII-D. If NEM is not crediting BTM DG of prosumer at a higher production rate $\pi^{-}$ than the expected wholesale LMP $\pi$, our aggregation method is profitable while competing with NEM to attain customers. In practice, $\pi^{-}\leq\mathbb{E}[\pi]\leq\pi^{+}$ happens since many states including California are setting the export rate $\pi^{-}$ at the avoided cost rate, which is, in many cases close to the expected LMP $\pi$, to mitigate cross-subsidies [33, 6, 4].
IV DERA Wholesale Market Participation
The virtual storage model is adopted by ISOs to enable DERAs’ participation in the wholesale market with bi-directional monetary and power flows [28, 29]. This means DERA can submit a combination of supply offers and demand bids, purchasing its aggregated consumption (as charging the virtual storage) and selling its aggregated production (as discharging).
IV-A Offer/Bid Curves of DERA in Energy Markets
As a virtual storage participant in the real-time energy market, the DERA is either self-scheduled or scheduled by ISO/RTO according to its bids and offers. This work focuses on developing price-quantity bid/offer curves that define DERA’s willingness to consume/produce. In a competitive market, such curves are the marginal cost of production and the marginal benefit of consumption derived from the optimal DERA decision in Theorem 1.
Let $Q$ be DERA’s aggregated quantity to buy (when $Q<0$) or sell (when $Q>0$) and $\pi$ be the wholesale market LMP. Let $G:=\sum_{n=1}^{N}g_{n}$ be the BTM DG aggregated by DERA. In a competitive market, a price-taking DERA participant bids truthfully with its aggregated supply function
$\begin{array}[]{l}Q=\mathscr{F}(\pi),~{}~{}\mathscr{F}(\pi):=G-\sum_{n=1}^{N}\mathbf{1}^{\top}{\bm{d}}^{*}_{n}(\pi),\end{array}$ (13)
where ${\bm{d}}^{*}_{n}$ is defined in (7). Note that the inverse of the DERA supply function $\mathscr{F}^{-1}(Q)$ defines the offer/bid curves of the DERA. For a quantity bid, the DERA forecasts the LMP and computes the optimal net production with (13). In contrast, for a price-quantity bid/offer curve $\mathscr{F}^{-1}(Q)$, the DERA avoids LMP forecasting, as the ISO clears the market using the submitted curve and ensures consistency between the LMP and the resulting dispatch. This is further explained in Lemma 1 below. Note also that the supply function depends on the aggregated BTM generation $G$, which is not known to the DERA at the time of the market auction. In practice, $G$ can be approximated by using historical data or $N\mathbb{E}(g_{n})$ via the Law of Large Numbers involving $N$ independent prosumers or via the Central Limit Theorem for independent and dependent random variables [34].
IV-B Market Efficiency with DERA Participation
We now establish that the DERA’s participation in the wholesale market achieves the same social welfare as that when all profit-maximizing prosumers participate in the wholesale market individually. We assume the wholesale market is competitive, where all participants are price takers with truthful bidding incentives. Prosumer notations in this section overlap with those in Sec. III, but subscripts are modified to include the transmission network bus index.
Consider a transmission network with $I$ buses. At each bus of the transmission network, we assume $M$ PoAs at the distribution network are connected and $N$ prosumers are aggregated by the proposed DERA model. Denote $U_{in}$ as the concave utility function for the $n$-th prosumer at the $i$-th transmission network bus. ${\bm{g}}_{n}:=(g_{in})_{i\in[I]}$ and ${\bm{d}}_{n}:=(d_{in})_{i\in[I]}$ are respectively the vectors of BTM DG generation and energy consumption for the prosumers. For simplicity, we ignore the number of energy-consuming devices for each prosumer, i.e., $K=1$, in this section. At each transmission bus, we sum up all load-serving entities and generators into one demand function and supply function. The load-serving entity at bus $i$ purchases electricity $e_{i}$ with a concave benefit function $B_{i}(e_{i})$. The generator at bus $i$ produces $p_{i}$ with a convex cost function $C_{i}(p_{i})$. Denote ${\bm{e}}:=(e_{i})_{i\in[I]}$, ${\bm{p}}:=(p_{i})_{i\in[I]}$. ${\bm{f}}\in\mathbb{R}^{L}$ is the line flow limit for $L$ branches of the transmission network. ${\bm{S}}\in\mathbb{R}^{L\times I}$ is the network parameter for DC power flow model.
Lemma 1 (Wholesale market clearing with DERA).
When prosumers participate in the wholesale market indirectly through the proposed DERA model with offer/bid curve (13), social welfare $\mbox{\sf SW}_{\mbox{\rm\tiny DERA}}$ is the optimal value of convex problem
$\underset{{\bm{D}},{\bm{p}},{\bm{e}}\geq 0}{\rm max}$ $~{}~{}\sum_{i=1}^{I}(\sum_{n=1}^{N}U_{in}(d_{in})+B_{i}(e_{i})-C_{i}(p_{i}))$ (14a)
[…]
$\lambda:$ $\sum_{i=1}^{I}p_{i}=\sum_{i=1}^{I}(\sum_{n=1}^{N}(d_{in}-g_{in})+e_{i}),$ (14b)
[…]
The sum of the DERA surplus and prosumers’ surpluses, denoted by $S_{\mbox{\tiny DERA}}$, can be computed by
$\begin{array}[]{l}S_{\mbox{\tiny DERA}}=\sum_{i=1}^{I}\sum_{n=1}^{N}(U_{in}(d^{\star}_{in})-\pi_{i}(d^{\star}_{in}-g_{in})),\end{array}$ (15)
where $d^{\star}_{in}$ is the optimal solution of (14), which equals (7).
[…]
As for the prosumer’s direct participation in the wholesale market, a price-taking prosumer $n$ and bus $i$ constructs her offer/bid curves by solving the following surplus maximization problem with the given LMP $\pi_{i}$:
$\begin{array}[]{l}\underset{d_{in}\in{\cal D}_{in}}{\rm max}~{}~{}U_{in}(d_{in})-\pi_{i}(d_{in}-g_{in}),\end{array}$ (16)
where ${\cal D}_{in}:=[\max\{\underline{d}_{in},g_{in}-\overline{c}_{in}\},\min\{\bar{d}_{in},g_{in}+\underline{c}_{in}\}]$. The access limits $\overline{c}_{in}$ and $\underline{c}_{in}$ represent the distribution network injection and withdrawal capacities allocated to each prosumer. [^8] Detail formulations are in (51)(52) of the appendix. These values are consistent with those computed in (7), ensuring a fair comparison. Under this setup, prosumers participating directly in the wholesale market face the same access constraints at each PoA as those in our proposed DERA model. By solving (16), we obtain the bid/offer curve for prosumer $n$ at bus $i$, i.e.,
$\begin{array}[]{l}\mathscr{S}_{in}(\pi_{i})=g_{in}-d^{*}_{in}(\pi_{i}),\end{array}$ (17)
where $d^{*}_{in}(\pi_{i})$ takes the same definition as that in (7).
Let $\mbox{\sf SW}_{\mbox{\rm\tiny Direct}}$ and $S_{\mbox{\rm\tiny PRO}}$ be, respectively, the optimal social welfare and prosumers’ surplus when all prosumers directly participate in the wholesale market. The following theorem is parallel to [12], although we employ different aggregation methods and consider distribution network access limits.
Theorem 2 (Market efficiency).
When all prosumers directly participate in the wholesale market, the market clearing result can be computed by (14), $\mbox{\sf SW}_{\mbox{\rm\tiny Direct}}=\mbox{\sf SW}_{\mbox{\rm\tiny DERA}}$, and $S_{\mbox{\rm\tiny PRO}}=S_{\mbox{\rm\tiny DERA}}.$
The sketch of proof is provided in Appendix VIII-F, which relies on the fact that the proposed DERA has its bidding curve (13) equal to the sum of the prosumer’s bidding curve in (17). From this, we can establish that the wholesale market clearing problem with the direct participation of all prosumers has the same market-clearing results as (14).
Although the proposed DER aggregation model only focuses on DERA’s profit maximization in the objective of (6), the competitive constraint (5) aligns the aggregated prosumer’s surplus maximization with DERA’s profit maximization. So, the proposed competitive DER aggregation has the incentive to maximize prosumers’ surpluses and get the maximum total surplus that can be split among DERA and its aggregated prosumers. Essentially, the DERA acts as an intermediary, enabling prosumers to indirectly participate in the wholesale market. As the DERA earns a profit for providing this service, each prosumer receives a lower surplus than they would under direct participation (illustrated in Fig. 2). This is justified, since individual prosumers lack the scale required for direct participation in the wholesale market.
V DERA-DSO coordination
All generation and consumption resources aggregated by DERA need to bypass the distribution network to participate in the wholesale market. The DERA aggregation presented in this work ensures that the aggregated DER at each distribution network PoA is bounded by access limits imposed through the distribution network access limit auction in [25]. A DERA submits a bid curve in this auction representing its willingness to acquire access at PoAs in the distribution system.
We assume that a DERA is a price taker in the access limit auction. Therefore, the bid-in demand curve for network access from the DERA at a particular PoA is the marginal benefit (profit) from having a DER aggregation under the PoA. The maximum expected profit of DERA is given by
$\varphi(\overline{{\bm{C}}},\underline{{\bm{C}}}):=\mathbb{E}_{{\bm{g}},\pi}[\Pi(\overline{{\bm{C}}},\underline{{\bm{C}}})],$ (18)
where $\Pi(\overline{{\bm{C}}},\underline{{\bm{C}}})$ is the maximum DERA profit computed from the optimal value of (6), given the realized renewable generations over all buses and the realized LMP. Note that when participating in the forward network access auction, both the BTM DG and LMP are random.
V-A DERA Benefit Function for Distribution Network Access
The following Proposition provides an expression for the benefit function of DERA, $\varphi(\overline{{\bm{C}}},\underline{{\bm{C}}})$, which can be used as the bid curve of access limits submitted to the auction in [25].
Denote ${\bm{h}}_{n}(x):=\sum_{k=1}^{K}h_{nk}(x)$ and two thresholds
$\begin{array}[]{l}\overline{q}_{m}:=\overline{C}_{m}+\sum_{n=1}^{{\cal N}_{m}}{\bm{h}}_{n}(\pi),~{}\underline{q}_{m}:=-\underline{C}_{m}+\sum_{n=1}^{{\cal N}_{m}}{\bm{h}}_{n}(\pi).\end{array}$
We have the following form for DERA’s benefit function.
Proposition 3 (Benefit function for network access).
With the DERA profit maximization (6), the expected DERA surplus is
$\varphi({\underline{{\bm{C}}}},{\overline{{\bm{C}}}})=\mathbb{E}\big{\{}\sum_{m=1}^{M}(\underline{\phi}_{m}({\underline{C}_{m}})+\overline{\phi}_{m}({\overline{C}_{m}}))+\sum_{n=1}^{N}(\varrho_{n}-{\cal K}_{n})\big{\}},$ (19)
$\underline{\phi}_{m}({\underline{C}_{m}}):=\big{(}\sum_{n=1}^{{\cal N}_{m}}U_{n}({\bm{h}}_{n}(\underline{\xi}_{m}))-\pi\underline{C}_{m}\big{)}\mathbbm{1}\{\sum_{n=1}^{{\cal N}_{m}}g_{n}\leq\underline{q}_{m}\},$
$\overline{\phi}_{m}({\overline{C}_{m}}):=\big{(}\sum_{n=1}^{{\cal N}_{m}}U_{n}({\bm{h}}_{n}(\overline{\xi}_{m}))+\pi\overline{C}_{m}\big{)}\mathbbm{1}\{\overline{q}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}g_{n}\},$
$\varrho_{n}:=\big{(}U_{n}({\bm{h}}_{n}(\pi))-\pi({\bm{h}}_{n}(\pi)-g_{n})\big{)}\mathbbm{1}\{\sum_{n=1}^{{\cal N}_{m}}g_{n}\in(\underline{q}_{m},\overline{q}_{m})\}.$
The sketch of proof is provided in Appendix VIII-A. Here, we have $U_{n}({\bm{h}}_{n}(x))=\sum_{k=1}^{K}U_{nk}(h_{nk}(x))$ from the additive property of the utility function. The optimal DERA surplus is decomposed into three terms: the surplus dependent on the withdraw access $\underline{\phi}_{m}({\underline{C}_{m}})$, the surplus dependent on the injection access $\overline{\phi}_{m}({\overline{C}})$, and the surplus independent of the network access. $\varphi({\underline{{\bm{C}}}},{\overline{{\bm{C}}}})$ is separable across injection and withdrawal access over $M$ PoAs. Therefore, DERA can bid separately for the distribution network accesses at different PoAs when coordinating with DSO. At PoA $m$ with less BTM DG, i.e., $\sum_{n=1}^{{\cal N}_{m}}g_{n}\leq\underline{q}_{m}$, DERA’s benefit depends on the withdrawal access. Conversely, if there is more BTM DG, i.e., $\overline{q}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}g_{n}$, DERA’s benefit depends on the injection access. Related simulations are in Sec. VI-D.
V-B Long-Run Equilibrium for Competitive DERA
In a long-run competitive industry, we explore how many DERA can survive. DERAs compete to attract customers, attain distribution network access, and participate in the wholesale market. We assume all DERAs adopt the competitive DER aggregation method in (6). The condition for a competitive long-run equilibrium [35, P193] has two components: (i) the marginal benefit of DERA equals the marginal cost of DSO for providing the distribution network access, and (ii) all DERAs have profits equal to zero, i.e., DERA’s profit in conducting aggregation equals DERA’s payment to acquire distribution network access. Related derivations and simulations are in Sec. VI-E and Appendix VIII-G.
VI Case Studies
We compared the expected surplus distribution of different DER aggregation methods under varying network access limits and BTM DG generations. Under the access limit allocation framework in [25], distribution network reliability concerns are resolved if DERAs obey allocated distribution network access limits. So the distribution network topology was ignored in the simulation. We also computed the benefit function of DERA to the distribution network access and empirically evaluated the long-run equilibrium of DERA with multi-interval aggregation. The DERA bidding curve to the wholesale market was simulated in our previous paper [1].
VI-A Parameter Settings
Denote utility function for the aggregated customer as
$U(x)=\begin{cases}\alpha x-\frac{\beta}{2}x^{2},&0\leq x\leq\frac{\alpha}{\beta}\\ \frac{\alpha^{2}}{2\beta},&x>\frac{\alpha}{\beta}\end{cases},$ (20)
where $\alpha=\$0.4/\mbox{kWh},\beta=\$0.1/(\mbox{kWh})^{2}$ [27]. Let the marginal utility $V^{-1}\in[\underline{d},\bar{d}]$ for the consumption limits. [^9] We simulated the case with prosumer device number $K=1$ for simplicity.
We used NEMa and NEMp to represent the DER aggregation under NEM when prosumers were active and passive, respectively. Passive customers are not responsive to the retail prices, but active customers will optimize their energy consumption given the retail price and the BTM DG generations. Based on PG&E residential rate, we set $\pi^{+}=\$0.3/\mbox{kWh}$ for the NEM. We assumed $\pi^{-}=\mathbb{E}[\pi]$ and the fixed cost of NEM was covered by extracting fixed payment from DERA, so we simulated with $\pi^{0}=\$0$. Gao-Alshehri-Birge (GAB) represented the two-part pricing in [12], which allowed customers to sell BTM DG to the DERA while purchasing energy from its incumbent utility company. Detailed models for NEMa, NEMp, and GAB are provided in Appendix VIII-B. Our DER aggregation method was simulated in Co.NEMa and Co.GAB, competitive with NEMa and GAB, respectively. For Co.NEMa, we set profit ratio $\zeta$ at the upper bound in the proof of Proposition 2. For Co.GAB, we set $\zeta=1.05$ to provide 5% more customer surplus than the GAB competitor.
We considered the randomness of LMP and BTM DG generation using data sources from CAISO [36] and Pecan Street Dataport [37], respectively. The LMP $\pi$ was modeled as a Gaussian random variable with a mean of \$0.05/kWh and a standard deviation (STD) of \$0.01/kWh. The BTM DG generation $g$ was modeled as a Gaussian random variable with a mean ranging from 1.1 kWh to 5.1 kWh and a standard deviation of 0.2 kWh, truncated at $(0,+\infty)$. We generated 10,000 random scenarios for both the LMP and BTM DG. At a given PoA, we evaluated the expected per-customer surplus metrics based on the sample means from these scenarios.
VI-B Performances with Unlimited Distribution Network Access
Four observations below were drawn when all aggregators received plenty of distribution network accesses.
First, Co.NEMa and Co.GAB were at the Pareto front in Fig. 2, achieving the maximum social surplus as if all prosumers directly participated in the wholesale market. This verified Theorem 2. Note that we computed the Pareto front by adding up surpluses of DERA and customers, omitting surpluses of other units. This was because we adopted the price taker assumption in the wholesale market, thus surpluses of other units stayed the same in different DERA models. The blue dot, named Direct, represented the ideal case that prosumers directly participated in the wholesale market with bidding curve (17). The green rectangular contained aggregation methods achieving less DERA surplus and customer surplus than Co.NEMa, thus dominated by our proposed competitive DER aggregation method. Similarly, the orange rectangular was dominated by its top right corner, Co.GAB. This was because our aggregation methods efficiently participated in the wholesale market with aggregated resources and scheduled the aggregated customers at a consumption level with a higher customer surplus. When the expected BTM DG increased from 1.1kW to 5.1kW, comparing the left and right panels in Fig. 2, we observed the expected social surplus, which was the sum of DERA and customer surpluses, increased, because more BTM DG was sold to the wholesale market.
Figure 2: Expected surplus distribution and market efficiency with 80% DG adoption rate. Each shaded rectangular is dominated by its top right corner. (From the left to the right, the expected DG increases from 1.1 to 5.1 kW.)
Second, customers had the highest expected surplus in Co.NEMa and Co.GAB shown by the top of Fig. 3. Passive customers in NEMp had the least surplus because its scheduling had no awareness of DG generation. Customer surpluses almost overlapped in all cases at a low DG adopter ratio with fewer producers, since most aggregation benefits came from BTM DG of producers. When the DG adopter ratio increased, the expected customer surplus increased in all cases.
Figure 3: Expected surplus distributions v.s. network access ratio. (Top: expected customer surplus; bottom: expected DERA surplus.)
Third, when DG adopter ratio or the DG generation was low, Co.NEMa and Co.GAB achieved the highest expected DERA surplus, as is shown in the bottom of Fig. 3. When the DG adopter ratio and DG generation were high, GAB achieved the highest DERA surplus because GAB only aggregated producers. [^10] GAB achieved the Pareto front when all prosumers were producers, e.g. DG adopter ratio equal 100% and $\mathbb{E}[g]=5.1$kW. Co.NEMa always had DERA profit no less than zero since we chose $\zeta$ based on Proposition 2. The DERA surpluses under NEMp and NEMa were identical, as setting $\pi^{-}=\pi$ ensures that aggregated DGs are compensated at the wholesale market price, eliminating any surplus gained by aggregating active prosumers’ DG production.
Fourth, since NEM provided a higher surplus to customers with BTM DG, DERAs must commensurately reduce their own profits and share them with the customers to remain competitive with NEM. Therefore, in most cases of Fig. 3, the expected DERA surplus decreased when the DG adopter ratio increased. However, GAB witnessed an increasing DERA surplus when the DG adopter ratio increased because GAB only aggregated producers.
VI-C Performances with Limited Distribution Network Access
Here, we set distribution network access limits for each prosumer by $\overline{C}=\underline{C}=8\delta$ kW and varied the network access ratio $\delta$ from 0 to 1 to analyze the influence of limited distribution network access. First, as is shown in Fig. 4, either Co.NEMa or Co.GAB achieved the highest customer surplus or DERA surplus under a limited network access ratio. Second, when the network access ratio increased, customer surplus increased in most cases except NEMp which passively controlled DG. Third, the DERA surplus in all cases increased when the network access ratio increased. This was intuitive because DERAs needed distribution network access to deliver the aggregated DER and participate in the wholesale market.
Figure 4: Expected surplus distributions v.s. network access ratio with 50% DG adoption rate. (Top: expected customer surplus; bottom: expected DERA surplus when $\mathbb{E}[g]=$ 1.1 kW and 5.1 kW, respectively.)
VI-D Benefit Function of DERA for Distribution Network Access
We computed the bid-in benefit function of the proposed DERA model, i.e., (19), with $\zeta=1.01$ and 50 prosumer aggregated at a certain PoA. DERA was competing with NEM, and prosumers were passive. Figure 5 shows the expected benefit $\varphi$ of the DERA as a function of injection and withdrawal access, under varying levels of expected BTM DG generations.
In Fig.5 (left), DERAs with lower expected DG generation exhibit higher benefits and submit higher bid prices for withdrawal access, as indicated by the steeper slope of the benefit function. This is because, with less BTM generation, DERAs rely more on electricity withdrawn from the network. In Fig.5 (right), the benefit function decreases with higher DG generation—a counterintuitive result. This occurs because NEM offers greater surplus to customers with higher DG output, forcing DERAs to reduce their profit margins and share more benefits with customers to remain competitive.
Figure 5: DERA Benefit function $\varphi$. (Left: withdrawal access $-\underline{C}$; right: injection access $\overline{C}$.)
VI-E Long-Run Competitive Equilibrium of DERAs
In the long-run competitive equilibrium analysis with multi-interval aggregation of DERs, we assumed 200 DERAs initially existed and computed the expected number of surviving DERAs in the long run. For simplicity, we assumed DERAs were homogeneous and had the same setting as Sec. VI-D. Prosumers had the same expected DG generation created from the 24-hour roof-top solar data in Pecan Street [37]. [^11] Detail DG trajectories and the long-run equilibrium results for single-interval aggregation are shown in Appendix VIII-G, providing intuitions about long-run equilibrium for multi-interval aggregation here. We multiplied the mean of 24-hour DG by ϵ1+subscriptitalic-$\epsilon_{1}\in\mathbb{R}_{+}$ to simulate different DG installation capacities and sampled 10,000 random DG scenarios. DERA submitted the benefit function, as in Fig. 5, to acquire hourly distribution network access. Same as [25], the DSO cost function for providing distribution network access was assumed to be the sum of quadratics, $J(x)=\frac{1}{2}bx^{2}+ax$ with $a=\$0.009/\mbox{kWh},b=\$0.0005/(\mbox{kWh})^{2}$ for both the injection and withdrawal access. We multipied DSO’s cost $J$ by ϵ2+subscriptitalic-$\epsilon_{2}\in\mathbb{R}_{+}$ to simulate different levels of DSO’s costs.
Two observations were drawn from results in Fig. 6. First, when the DG capacity ratio $\epsilon_{1}$ was about 0.4-1.4, all initial 200 DERAs survived because DERAs can internally balance customer demands with its aggregated DG, thus relying and paying less to the network access. This was validated by the yellow dot curve from Fig. 6 (right), which required almost zero network access over 24 hours. Second, when the DG capacity ratio decreased from 0.4 to 0 in Fig. 6 (left), the number of surviving DERA decreased. In this case, DG was lower than the aggregated customers’ consumption, and not all DERAs can survive when competing and paying for the network withdrawal access over 24 hours, shown by the blue solid curve of Fig. 6 (right). In the green dash curve of Fig. 6 (left), DSO’s cost for providing network access was lower, so more DERAs survived than other curves. Similar reasons applied when DG capacity ratio increased beyond 1.4.
Figure 6: Long-run competitive equilibrium for multi-interval aggregation. (Left: expected number of surviving DERA v.s. DG installation capacity ratio $\epsilon_{1}$; right: expected distribution network net injection access of DERA over 24-hour, whose negativity represents withdrawal access.)
VII Conclusions
A major challenge to realizing the goals of FERC Order $2$ is enabling profit-maximizing DERAs to effectively compete with the retail programs offered by distribution utilities. To this end, this paper considers the competitive DER aggregation of a profit-seeking DERA in the wholesale electricity market. As a wholesale market participant, DERA can both inject and withdraw power from the wholesale market. It is shown that the proposed DERA model maximizes its profit while providing competitive services to its customers with higher surpluses than those offered by the distribution utilities. We also establish that the resulting social welfare from DERA’s participation on behalf of its prosumers is the same as that gained by the direct participation of price-taking prosumers, making the proposed DERA aggregation model optimal in achieving wholesale market efficiency. Additionally, we derive two significant optimal price-quantity bids of DERA, of which one is submitted to the wholesale market, and the other to the distribution network access allocation mechanism [25].
An open issue of the proposed aggregation solution is that the payment functions for prosumers are nonlinear and non-uniform. Although each customer is guaranteed to be better off than the competing scheme, two customers producing the same amount may be paid and compensated differently. In other words, the total charge/credits depend not only on the quantity but also on the flexibilities of the demand and constraints imposed by the prosumer. Note that a profit-seeking DERA participating in the wholesale electricity market is not subject to the same regulation as a regulated utility. Such non-uniform pricing is acceptable and has also been proposed in the form of non-uniform fixed charges [12, 6].
Acknowledgement
The authors are grateful for the many discussions with Dr. Tongxin Zheng, Dr. Mingguo Hong, and other colleagues from ISO New England. We also benefited from helpful comments and critiques from Dr. Jay Liu and Dr. Hong Chen from PJM.